OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3).

OpenStudy (anonymous):

y=(3/2)x+4 = perpendicular= y=(-2/3)x+4

OpenStudy (anonymous):

maybe (2/3)x+3

OpenStudy (anonymous):

yo bro can u help

OpenStudy (anonymous):

y=-(2/3)x+b 3=-(2/3)(-6)+b 3=4+b -4 -4 b=-1 so y=-(2/3)x-1